5.The Locus of points at a distance of 3 from the point (0, 0) is given by the equation…
a.x^2+y^2=−9

b.x^2-y^2=9

c.x^2+y^2=9

d.−x^2+y^2=−9

6.If A is (3, 6) and B is (4, 8). Find the coordinate of the midpoint of AB.
a.[3 1/2,−7]
b.[−3 1/2,7]
c.[−3 1/2,−7]
d.[3 1/2,7]
1
Expert's answer
2019-04-03T10:10:06-0400

Answer:

5) c

6) d


Explanation:

5)In other words we need choose circle equation with center (0, 0) and radius 3:

"(x-0)^2 + (y-0)^2 = 3^2"


6) the coordinates of midpoint:

"(\frac{x_a + x_b}{2}, \frac{y_a + y_b}{2}) = (\frac{3+4}{2}, \frac{6+8}{2}) = (3.5, 7)"